OpenStudy (anonymous):

x^3-2x^2+x-2=0 what does x =

OpenStudy (anonymous):

x^3-2x^2+x-2=0 =>x^2(x-2)+1(x-2) =0 =>(x^2-1)(x-2)= 0 => (x-1)(x+1)(x-2)= 0 x = 1,-1 or 2